Find the rate of return for the following cash flow. Year Cash Flow -$25,000 15.000 -4.000 8.000 | 15,000 14
Answer #1

To calculate rate of return (ROR) in the problem we have followed trial and error method

Present Value factor (PVF) = (1+i)-N

Years(N)

Cash flow

PVF@12%

i=12%

PV of Cashflow

PVF@13%

I=13%

PV of Cashflow

0

-25000

1

-25000

1

-25000

1

15000

0.8928571

13392.8571

0.8849558

13274.3363

2

-4000

0.7971939

-3188.7755

0.7831467

-3132.5867

3

8000

0.7117802

5694.2420

0.6930502

5544.4013

4

15000

0.6355181

9532.7712

0.6133187

9199.7809

NPV at 12%=

431.0948

NPV at 25%=

-114.0682

Rate of return (ROR) is the interest rate at which the NPV=0, to estimate it we have followed trial and error method, first we have taken the PV factor = 12%, where the NPV =$431.0948, which is positive and then we have taken the PV factor =13%, where the NPV = -$114.0682which is negative. So the Rate of return must fall in between 12% to 13%.

Therefore

R1 = 12%                             NPV1=$431.0948

R2= 13%                               NPV2=-$114.0682

Using linear interpolation method of estimation of Rate of Return

ROR= R1 + (R2-R1)X[NPV1/(NPV1-NPV2)]

ROR = 12 + (13-12)[( 431.0948)/ {431.0948-(-114.0682)}]

ROR = 12 + (1)[( 431.0948)/ (431.0948+114.0682)

ROR = 12 + (1)[( 431.0948)/ (545.163) =12 +0.7907    or Approx 12.79%

The annual rate of return is approx 12.79%